# Fast, robust and laser-free universal entangling gates for trapped-ion quantum computing

### v2

@@ Appendix C RF amplitude calibration
The gate scheme requires a well-calibrated Rabi frequency $\Omega_{1}^{\text{Amp}}$, matching the chosen parameter $\Omega_{1}^{\text{Phase}}$, explained in Eq. (5). Experimentally, we calibrate the physical Rabi frequency $\Omega_{1}^{\text{Amp}}$ by well known Rabi measurements. We address each qubit by a not phase-modulated on-resonant RF field near $2\pi\times 12.6$ GHz. We measure the excitation probability for different interaction times, and fit the data using a sinusoidal function, thus extracting the Rabi frequency $\Omega_{1}^{\text{Amp}}$. By accordingly changing the RF amplitude, we calibrate $\Omega_{1}^{\text{Amp}}\approx\Omega_{1}^{\text{Phase}}$.
In order to efficiently carry out this calibration, we also measure the nonlinear dependence between the amplitude of the RF signal and the observed Rabi frequency. Addressing one qubit transition on resonance using the RF signal generated with an AWG, we measure Rabi oscillations with frequencies up to $2\pi\times 200$ kHz. We show the measured frequencies in Fig. 7. Saturation sets in for RF amplitudes at about half the maximum amplitude. For desired Rabi frequencies close to the trap frequency, we operate in this nonlinear regime which can also affects the usage of amplitude-modulated fields.
Figure 7: For one ion, we measure the on-resonant Rabi frequency $\Omega$ as a function of the amplitude $A$ of the RF signal. The measurement accuracy is better than $2\pi\times 0.1$ kHz. The non-linearity is due to saturation of the RF amplifiers.
@@ Appendix D Tomography
where the coefficients $\lambda_{i}$ are the expectation values of the operators shown in Tab. 1. Following [58], we reconstruct the density matrix for a two-qubit system by measuring the expectation values $\langle\sigma_{i}\otimes\sigma_{j}\rangle,i,j=0,1,2,3$, where $\sigma_{i}$ runs over the set of Pauli matrices $\mathds{1},\sigma_{x},\sigma_{y},\sigma_{z}$. In the experiment described here, only the $\sigma_{z}$ eigenvalue can be measured directly by a projective measurement detecting resonance fluorescence near $369$ nm. Therefore, in the two-qubit system investigated here, we can only directly measure the observables
@@
The corresponding expectation values are calculated using the experimental probabilities as,
@@
Measuring the expectation values of $\sigma_{x}$ and $\sigma_{y}$is realized by mapping the quantum state onto the eigenvector of $\sigma_{z}$, by applying single qubit rotations $R(\theta,\phi)$. We apply nine different sets of qubit rotations, shown in Tab. 1, to extract all 16 expectation values. $\lambda^{(k)}_{i},$ is the expectation value of the observable $O_{i}$ after the qubit rotation $(k)$, shown in the $k$-th row of Tab. 1, is applied. The reconstructed density matrix then is given by
@@
k Ion 1 Ion 2 Exp. values
1 - - $\langle\sigma_{z}\otimes\mathds{1}\rangle$ $\langle\mathds{1}\otimes\sigma_{z}\rangle$ $\langle\sigma_{z}\otimes\sigma_{z}\rangle$
2 $23\pi/$ 2 - $\langle\sigma_{x}\otimes\mathds{1}\rangle$ $\langle\sigma_{x}\otimes\sigma_{z}\rangle$
3 $\pi$ $\langle\sigma_{y}\otimes\mathds{1}\rangle$ $\langle\sigma_{y}\otimes\sigma_{z}\rangle$
4 - $23\pi/$ 2 - $\langle\mathds{1}\otimes\sigma_{x}\rangle$ $\langle\sigma_{z}\otimes\sigma_{x}\rangle$
5 - $\pi$ $\langle\mathds{1}\otimes\sigma_{y}\rangle$ $\langle\sigma_{z}\otimes\sigma_{y}\rangle$
6 $23\pi/$ 2 $23\pi/$ 2 - - $\langle\sigma_{x}\otimes\sigma_{x}\rangle$
7 $23\pi/$ 2 $\pi$ $\langle\sigma_{x}\otimes\sigma_{y}\rangle$
8 $\pi$ $23\pi/$ 2 - - $\langle\sigma_{y}\otimes\sigma_{x}\rangle$
9 $\pi$ $\pi$ $\langle\sigma_{y}\otimes\sigma_{y}\rangle$
Table 1: The density matrix of a two-qubit system is reconstructed by mapping the quantum state to the $\sigma_{z}$ basis. This is realized applying either one of the transformations $k$, which correspond to single-qubit rotations $R(\pi/2,\phi)$. Subsequently, the state is read out in the $\sigma_{z}$ basis to reconstruct expectation values.
Assuming standard errors and using Gaussian error propagation, the variance of the real and imaginary parts of $\rho$ are obtained.
@@ Appendix E Negativity
In this work, we use negativity $\mathcal{N}$ to quantify the degree of entanglement. $\mathcal{N}$ is a state-indepedent measure of entanglement, which is defined as the absolute value of the sum of the negative eigenvalues of $\rho^{\Gamma_{A}}$ [59]. Here, $\rho^{\Gamma_{A}}$ is the partial transpose of the two-qubit density matrix with respect to the subset of the first qubit. For a maximally entangled state, the negativity is equal to 0.5. The measure is meaningful, since a mixed two-qubit state is entangled if and only if its partial transpose has a negative eigenvalue. The statistical error of this measure is calculated by generating a set of density matrices based on a Gaussian probability distribution using the entries of the measured density matrix as mean and the statistical error as width of the distribution (see Fig. 8). Sampling the negativity of all generated density matrices, give the standard deviation of the measured density matrix.
Figure 8: Numerically obtained probability distribution of the negativity for the |$\ket{\Phi^{+}}$ (a) and |$\ket{\Psi^{+}}$ (b) Bell state. The histogram is synthesized by sampling a set of density matrices based on a Gaussian probability distribution using the entries of the measured density matrix as mean and the statistical error as width of the distribution. The orange line shows the mean of the probability distribution, and the yellow line indicates the directly calculated value of the original measured density matrix. The dashed line shows the optimal value for a pure, maximally entangled state.
@@ Appendix F Gate evolution in the computational basis
Figure 9: The gate’s time evolution is measured in terms of the excitation probability in the bare state basis. The ions are initialized in the eigenstate of the dressed basis and the gate field is applied for duration $t$ Then, the ions’ states are detected. Each data point represents the result of 100 repetitions of the experiment. The orange line indicates the gate time used to implement the entangling gates for the Bell state |$\ket{\Psi^{+}}$, and the blue dashed line marks the gate time for the Bell state |$\ket{\Phi^{+}}$. The solid black line shows the simulated gate evolution, using parameters that best match the experimental data: $\nu=2\pi\times 97.85$ kHz, $\Omega_{1}=2\pi\times 94.83$ kHz, and $\Omega_{2}=23\times\eta\nu$.
To investigate the full gate evolution in terms of the excitation probability in bare state basis, we apply a resonant $\pi/2$-pulse to both ions, bringing them into a superposition state. Then, the phase modulated driving field interacts with the ions for time $t$. Fig. 9 shows the two-qubit product state excitation probabilities $P_{|ij\rangle}$ ($i$, $j$ denote qubit states.) The left column presents the initial stage of the time evolution, while the right column presents the time evolution around the entangling gate time. The experimental results are overlapped with simulation results agreeing well with the experimental results by using the parameters $\nu=2\pi\times 97.85$ kHz, $\Omega_{1}=2\pi\times 94.83$ kHz, and $\Omega_{2}=23\times\eta\nu$ for the simulation. The dashed blue (orange) vertical line illustrates the selected gate time at which the excitation probability matches the |$\ket{\Phi^{+}}$ (|$\ket{\Psi^{+}}$) Bell state.
@@ Appendix G Detection correction
The qubit readout is realized by collecting resonance fluorescence using a global laser beam close to $369$ nm. The fluorescence of each ion is imaged onto an EMCCD camera, producing a spatially resolved image of the ion chain. Using a two-threshold method to analyze the collected fluorescence, we distinguish the qubit’s logical states |$\ket{0}$ and |$\ket{1}$ [68]. For the experiments reported here, the product state probabilities $P_{i},\leavevmode\nobreak\ i=1,2,3,4$ of the computational basis states $\{1:\ket{00},2:\ket{01},3:\ket{10},4:\ket{11}\}$ need to be reconstructed from the measured relative frequencies $\widetilde{P_{i}}$ assigned using double threshold detection. Due to imperfect spatial separation of the ions’ fluorescence, the readout may be assigned wrongly. To account for this detection error, we carry out a correction of the detected excitation probabilities of the product states. To this end, we prepare all possible computational basis states by first optically pumping into state |$\ket{00}$, then applying a Rapid Adiabatic Passage (RAP)[69] pulse, specific for the state to prepare, and finally read out the ions’ internal state. In this way, the probabilities of wrong assignments of readouts can be inferred. The state vector
$\vec{P}$ $=(P_{1},P_{2},P_{3},P_{4})\;,$ (42)
therefore, describes the excitation probability of all computational basis states. A linear map, $M$ between the probability vector $\vec{P}$ and the measured relative frequencies $\vec{\widetilde{P}}=M\vec{P}$ can be found. Here, $M$ is a real-valued $4\times 4$ matrix, and the elements of $M$ are given by the measured relative frequencies, i.e $M_{ji}=\widetilde{P}_{i},\leavevmode\nobreak\ j=1,2,3,4$ corresponding to the prepared states. This linear map, therefore, contains all possible wrong assignments. Applying the inverse matrix $\vec{P}=M^{-1}$ allows for reconstructing the excitation probabilities of the ions’ state, compensating for detection errors: $\vec{P}=M^{-1}\vec{\widetilde{P}}$. For the two-qubit system, we show as an example the corresponding detection matrix in Fig. 10. Assuming that all statistical errors follow a normal distribution, the error of the readout correction can be propagated following the descriptions in [70]. The standard deviation of the inverse matrix elements is
@@
Using Gaussian error propagation, the standard errors of the reconstructed state is
@@
Since the correction of readout error is not guaranteed to be unitary, using this correction for detection errors might result in nonphysical probabilities exceeding the interval $[0,1]$. In further analysis, we therefore truncate nonphysical quantities. For the |$\ket{\Psi^{+}}$(|$\ket{\Phi^{+}}$) Bell state, the reconstructed density matrix has one negative eigenvalue of -0.04(-0.05).

### v3

@@ Appendix C RF AMPLITUDE CALIBRATION
The gate scheme requires a well-calibrated Rabi frequency $\Omega_{1}^{\text{amp}}$, matching the chosen parameter $\Omega_{1}^{\text{phase}}$, explained in Eq. (5). Experimentally, we calibrate the physical Rabi frequency $\Omega_{1}^{\text{amp}}$ by well-known Rabi measurements. We address each qubit by an on-resonant rf field near $2\pi\times 12.6$ GHz that is not phase modulated. We measure the excitation probability for different interaction times and fit the data using a sinusoidal function, thus extracting the Rabi frequency $\Omega_{1}^{\text{amp}}$. By accordingly changing the rf amplitude, we calibrate $\Omega_{1}^{\text{amp}}\approx\Omega_{1}^{\text{phase}}$.
In order to efficiently carry out this calibration, we also measure the nonlinear dependence between the amplitude of the rf signal and the observed Rabi frequency. Addressing one qubit transition on resonance using the rf signal generated with an AWG, we measure Rabi oscillations with frequencies up to $2\pi\times 200$ kHz. We show the measured frequencies in Fig. 7. Saturation sets in for rf amplitudes at about half the maximum amplitude. For desired Rabi frequencies close to the trap frequency, we operate in this nonlinear regime, which can also affect the usage of amplitude-modulated fields.
Figure 7: For one ion, we measure the on-resonant Rabi frequency $\Omega$ as a function of the amplitude $A$ of the rf signal. The measurement accuracy is better than $2\pi\times 0.1$ kHz. The nonlinearity is due to saturation of the rf amplifiers.
@@ Appendix D TOMOGRAPHY
where the coefficients $\lambda_{i}$ are the expectation values of the operators shown in Table 1. Following Ref. [58], we reconstruct the density matrix for a two-qubit system by measuring the expectation values $\langle\sigma_{i}\otimes\sigma_{j}\rangle,i,j=0,1,2,3$, where $\sigma_{i}$ runs over the set of Pauli matrices $\mathds{1},\sigma_{x},\sigma_{y},\sigma_{z}$. In the experiment described here, only the $\sigma_{z}$ eigenvalue can be measured directly by a projective measurement detecting resonance fluorescence near $369$ nm. Therefore, in the two-qubit system investigated here, we can only directly measure the observables
@@
The corresponding expectation values are calculated using the experimental probabilities
@@
Measuring the expectation values of $\sigma_{x}$ and $\sigma_{y}$is performed by mapping the quantum state onto the eigenvector of $\sigma_{z}$, applying single-qubit rotations $R(\theta,\phi)$. We apply nine different sets of qubit rotations, shown in Table 1, to extract all 16 expectation values. Note that $\lambda^{(k)}_{i},$ is the expectation value of the observable $O_{i}$ after the qubit rotation $(k)$, shown in the $k$th row of Table 1, is applied. The reconstructed density matrix is then given by
@@
k $\phi$ Ion 1 $\phi$ Ion 2 Expectation values
1 $\cdots$ $\cdots$ $\langle\sigma_{z}\otimes\mathds{1}\rangle$ $\langle\mathds{1}\otimes\sigma_{z}\rangle$ $\langle\sigma_{z}\otimes\sigma_{z}\rangle$
2 $23\pi/$ 2 $\cdots$ $\langle\sigma_{x}\otimes\mathds{1}\rangle$cdots $\langle\sigma_{x}\otimes\sigma_{z}\rangle$
3 $\pi$cdots $\langle\sigma_{y}\otimes\mathds{1}\rangle$cdots $\langle\sigma_{y}\otimes\sigma_{z}\rangle$
4 $\cdots$ $23\pi/$ 2 $\cdots$ $\langle\mathds{1}\otimes\sigma_{x}\rangle$ $\langle\sigma_{z}\otimes\sigma_{x}\rangle$
5 $\cdots$ $\pi$cdots $\langle\mathds{1}\otimes\sigma_{y}\rangle$ $\langle\sigma_{z}\otimes\sigma_{y}\rangle$
6 $23\pi/$ 2 $23\pi/$ 2 $\cdots$ $\cdots$ $\langle\sigma_{x}\otimes\sigma_{x}\rangle$
7 $23\pi/$ 2 $\pi$cdots $\cdots$ $\langle\sigma_{x}\otimes\sigma_{y}\rangle$
8 $\pi$ $23\pi/$ 2 $\cdots$ $\cdots$ $\langle\sigma_{y}\otimes\sigma_{x}\rangle$
9 $\pi$ $\pi$cdots $\cdots$ $\langle\sigma_{y}\otimes\sigma_{y}\rangle$
Table 1: Set of transformations $k$ (single-qubit rotations $R(\pi/2,\phi)$ that are experimentally applied to map two-qubit states to the $\sigma_{z}$ basis. Subsequently, the state is read out in the $\sigma_{z}$ basis. Using the measured expectation values the density matrix of the two-qubit system is reconstructed
Assuming standard errors and using Gaussian error propagation, the variance of the real and imaginary parts of $\rho$ is obtained.
@@ Appendix E NEGATIVITY
In this work, we use negativity $\mathcal{N}$ to quantify the degree of entanglement. Note that $\mathcal{N}$ is a state-independent measure of entanglement, which is defined as the absolute value of the sum of the negative eigenvalues of $\rho^{\Gamma_{A}}$ [59]. Here, $\rho^{\Gamma_{A}}$ is the partial transpose of the two-qubit density matrix with respect to the subset of the first qubit. For a maximally entangled state, the negativity is equal to 0.5. The measure is meaningful since a mixed two-qubit state is entangled if and only if its partial transpose has a negative eigenvalue. The statistical error of this measure is calculated by generating a set of density matrices based on a Gaussian probability distribution using the entries of the measured density matrix as the mean and the statistical error as the width of the distribution (see Fig. 8). Sampling the negativity of all generated density matrices, gives the standard deviation of the measured density matrix.
Figure 8: Numerically obtained probability distribution of the negativity for the |$\ket{\Phi^{+}}$ (a) and |$\ket{\Psi^{+}}$ (b) Bell states. The histogram is synthesized by sampling a set of density matrices based on a Gaussian probability distribution using the entries of the measured density matrix as the mean and the statistical error as the width of the distribution. The orange line shows the mean of the probability distribution, and the yellow line indicates the directly calculated value of the original measured density matrix. The dashed line shows the optimal value for a pure, maximally entangled state.
@@ Appendix F GATE EVOLUTION IN THE COMPUTATIONAL BASIS
Figure 9: Gate’s time evolution measured in terms of the excitation probability in the bare state basis. The ions are initialized in the eigenstate of the dressed basis, and the gate field is applied for duration $t$. Then, the ions’ states are detected. Each data point represents the result of 100 repetitions of the experiment. The orange line indicates the gate time used to implement the entangling gates for the Bell state |$\ket{\Psi^{+}}$, and the blue dashed line marks the gate time for the Bell state |$\ket{\Phi^{+}}$. The solid black line shows the simulated gate evolution, using parameters that best match the experimental data: $\nu=2\pi\times 97.85$ kHz, $\Omega_{1}=2\pi\times 94.83$ kHz, and $\Omega_{2}=23\times\eta\nu$.
To investigate the full gate evolution in terms of the excitation probability in the bare state basis, we apply a resonant $\pi/2$ pulse to both ions, bringing them into a superposition state. Then, the phase-modulated driving field interacts with the ions for time $t$. Figure 9 shows the two-qubit product state excitation probabilities $P_{|ij\rangle}$ ($i$, $j$ denote qubit states). The left column presents the initial stage of the time evolution, while the right column presents the time evolution around the entangling gate time. The experimental results are overlapped with simulation results agreeing well with the experimental results by using the parameters $\nu=2\pi\times 97.85$ kHz, $\Omega_{1}=2\pi\times 94.83$ kHz, and $\Omega_{2}=23\times\eta\nu$ for the simulation. The dashed blue (orange) vertical line illustrates the selected gate time at which the excitation probability matches the |$\ket{\Phi^{+}}$ (|$\ket{\Psi^{+}}$) Bell state.
@@ Appendix G DETECTION CORRECTION
The qubit readout is realized by collecting resonance fluorescence using a global laser beam close to $369$ nm. The fluorescence of each ion is imaged onto an EMCCD camera, producing a spatially resolved image of the ion chain. Using a two-threshold method to analyze the collected fluorescence, we distinguish the qubit’s logical states |$\ket{0}$ and |$\ket{1}$ [68]. For the experiments reported here, the product state probabilities $P_{i},\leavevmode\nobreak\ i=1,2,3,4$ of the computational basis states $\{1:\ket{00},2:\ket{01},3:\ket{10},4:\ket{11}\}$ need to be reconstructed from the measured relative frequencies $\widetilde{P_{i}}$ assigned using double threshold detection. Because of the imperfect spatial separation of the ions’ fluorescence, the readout may be wrongly assigned. To account for this detection error, we carry out a correction of the detected excitation probabilities of the product states. To this end, we prepare all possible computational basis states by first optically pumping into state |$\ket{00}$, then applying a rapid adiabatic passage (RAP)[69] pulse (specific for the state to prepare), and finally reading out the ions’ internal state. In this way, the probabilities of wrong assignments of readouts can be inferred. Therefore, the state vector
$\vec{P}$ $=(P_{1},P_{2},P_{3},P_{4})$ (42)
describes the excitation probability of all computational basis states. A linear map, $M$ between the probability vector $\vec{P}$ and the measured relative frequencies $\vec{\widetilde{P}}=M\vec{P}$ can be found. Here, $M$ is a real-valued $4\times 4$ matrix, and the elements of $M$ are given by the measured relative frequencies, i.e., $M_{ji}=\widetilde{P}_{i},\leavevmode\nobreak\ j=1,2,3,4$, corresponding to the prepared states. This linear map therefore contains all possible wrong assignments.
Applying the inverse matrix $\vec{P}=M^{-1}$ allows for reconstructing the excitation probabilities of the ions’ state, compensating for detection errors: $\vec{P}=M^{-1}\vec{\widetilde{P}}$. For the two-qubit system, we show as an example the corresponding detection matrix in Fig. 10. Assuming that all statistical errors follow a normal distribution, the error of the readout correction can be propagated following the descriptions in Ref. [70]. The standard deviation of the inverse matrix elements is
@@
Using Gaussian error propagation, the standard error of the reconstructed state is
@@
Since the correction of the readout error is not guaranteed to be unitary, using this correction for detection errors might result in nonphysical probabilities exceeding the interval $[0,1]$. In further analysis, we therefore truncate nonphysical quantities. For the |$\ket{\Psi^{+}}$(|$\ket{\Phi^{+}}$) Bell state, the reconstructed density matrix has one negative eigenvalue of –0.04(–0.05).